# QuantumBind-RBFE: Accurate Relative Binding Free Energy Calculations Using Neural Network Potentials

### v1

@@ Abstract
Accurate prediction of protein-ligand binding affinities is crucial in drug discovery, particularly during hit-to-lead and lead optimization phases, however, limitations in ligand force fields continue to impact prediction accuracy. In this work, we validate relative binding free energy (RBFE) accuracy using neural network potentials (NNPs) for the ligands. We utilize a novel NNP model, AceForce 1.0, based on the TensorNet architecture for small molecules that broadens the applicability to diverse drug-like compounds, including all important chemical elements and supporting charged molecules. Using established benchmarks, we show overall improved accuracy and correlation in binding affinity predictions compared with GAFF2 for molecular mechanics and ANI2-x for NNPs. Slightly less accuracy but comparable correlations with OPLS4. We also show that we can run the NNP simulations at 2 fs timestep, at least two times larger than previous NNP models, providing significant speed gains. The results show promise for further evolutions of free energy calculations using NNPs while demonstrating its practical use already with the current generation. The code and NNP model are publicly available for research use.
@@ 1 Introduction
Recently, an alternative approach has emerged based on neural network potentials (NNPs) to use fast neural network function approximation of the quantum mechanic’s energy surface12. Based on these architectures, a few methods have been made available which we will consider in this work. MACE-based models13 are very accurate but computationally expensive. This is not a problem in material science where the alternative is very expensive quantum chemistry methods. For biology-oriented applications, ANI-models14 are much faster but at the expenses of accuracy, still providing a very good trade-off. AIMNet215 extends ANI with a different architecture to more elements and molecules, still adding on computational costs. The equivariant transform (ET) architecture16 and TensorNet17 aimed to provide state-of-the-art accuracy without sacrificing computational efficiency, being taylored specifically for drug discovery. These two are integrated in the TorchMD-Net software framework 18, 16, 19 and a new NNP model has been made available, AceForce20. Molecular dynamics (MD) simulations based on NNPs promise accurate simulations at reasonable costs compared with density functional theory (DFT) methods, but do represent an increase in computational costs compared to traditional molecular mechanics, yet their future is very promising 21.
@@
Recently, we explored the integration of NPs into RBFE calculations with promising results.22 Specifically, we used ANI-2x14 with good results. However, ANI-2x cannot handle charged species and has a restricted range of supported atom elements. These limitations restrict the applicability across the diverse chemical space encountered in drug discovery. In this work, we extend our previous work22 by testing QuantumBind-RBFE, our NNP/MM approach23 for RBFE calculations, using AceForce. The AceForce 1.020 model supports a broad range of atom elements, including charged molecules, and is tailored for accurate and efficient RBFE calculations across diverse molecular systems. To validate the performance of RBFE calculations using NNPs, we conducted a comprehensive benchmarking study using established datasets, including both charged and neutral ligands. Our results show that we can improve the accuracy of RBFE calculations compared to traditional force fields and our previous test with ANI-2x, reaching state-of-the-art correlations.
@@ 2.1 RBFE calculation within the NNP/MM scheme
Calculations were performed using an NNP/MM approach, which combines neural network potentials (NNP) for high-accuracy modeling of ligand interactions and molecular mechanics (MM) for the remainder of the system.23 This hybrid method allows the ligand to be simulated with the more accurate neural network potential, while the surrounding protein environment is treated with classical molecular mechanics for speed.(Figure 1) This setup is particularly advantageous for capturing the finer details of intra-molecular ligand interactions that are typically challenging for traditional MM force fields, like ligand strain.
In NNP/MM, the system is divided into NNP and MM regions, akin to the partitioning used in QM/MM simulations. The total potential energy ($V$) is calculated as:
@@
where $V_{\text{NNP}}$ and $V_{\text{MM}}$ represent the potential energies for the NNP and MM regions, respectively, and $V_{\text{NNP-MM}}$ accounts for the coupling between these regions. The NNP potential is computed based on the atomic positions and charges within the NNP region, ensuring smooth interaction with the MM-treated environment. 23
@@
We utilized the Alchemical Transfer Method (ATM) to perform the RBFE calculations.24 This methodology has been previously validated across multiple benchmarks,25, 26 demonstrating its reliability for accurate binding free energy predictions. Notably, ATM was also employed in our recent work validating its use with NNPs, specifically ANI2x.22 In this study, we extend the previous work substantially.
@@ 2.2 Dataset selection
We selected a widely recognized public benchmark dataset that serves as a reference for RBFE studies. The BACE, CDK2, JNK1, MCL1, P38, THROMBIN, and TYK2 targets are part of Wang et al.’s dataset, commonly known as the ”JACS dataset” or ”Schrödinger dataset” 27. We had to omit PTP1B from the evaluation study since all ligands in the series have charge -2 and AceForce, for now, can only compute ligands with charges -1,0 and +1. Compared to our previous work, we can now perform the benchmark in completion. This dataset provides a total of 7 protein targets, 179 ligands, and 280 edges. Comparisons will be made with GAFF2 and FEP+ with the OPLS4 force field28.
@@ 2.3 NNP model
AceForce 1.0 is the first version of a new family of potentials 20. It uses TensorNet 1-layer 17 trained on Acellera29’s internal proprietary dataset of forces and energies computed the wB97M-V/def2-tzvppd level of theory. This work is the first test of this model for accuracy in RBFE calculations. A basic evaluation of the model on the Seller’s torsion scan 30 benchmark shows promising performance (Figure 2). For all molecules in the test, we performed the full torsion scans, minimizing the geometry with constrained torsion angles using the forces and energy from each calculation method. For comparison, we did the same procedure with ANI-2x14 from TorchANI31, xTB32, AIMNet215, and MACE-OFF13. We used GeomeTRIC33 for the constrained optimizations. The MAE between the curves for each molecule is shown in Figure 2. The AceForce model is among the best performing and comparable to MACE but at a much faster computational speed. Also, AceForce is using a faster TensorNet 1-layer model, instead to the 2-layer model generally used in 17 for maximum accuracy.
Figure 2: Seller’s torsion scan benchmark, comparing torsion accuracy against a high level quantum chemistry approximation.
@@ 2.4 QuantumBind-RBFE Calculation details
The workflow in this project follows the same procedure as in our previous work22. Ligands were parameterized with GAFF26, 7 with RESP charges, which are calculated with the OpenFF-Recharge package.34 The topologies were generated using the parameterize35 tool. Next, complex systems were prepared with HTMD,36 and atom indexes for ligand alignment were performed, selecting three reference atoms for each ligand. For more information on the selection of these atoms please check previous publications.24, 25.
The energy minimization, thermalization, and equilibration steps followed the procedures described in our previous work.25. RBFE simulations were run in triplicate for each edge for an ensemble of 70 ns per replica, a similar amount of simulation time to our previous22 and similar37 studies. We used the Amber ff14SB parameters38, 39 as well as the TIP3P water model. Classical RBFE simulations were run at a $4$ fs timestep while the NNP/MM runs were run at $2$ fs timestep. Previously, we already evaluated how the accuracy difference between 1fs and 4fs runs is minimal for MM runs.22 In the results section we discuss the reasoning behind selecting such a timestep for the NNP/MM runs. All calculations were run on GPUGRID.net40. The parallel replica exchange molecular dynamics simulations were conducted using QuantumBind-RBFE our implementation of ATM, HTMD36 and the ATMForce potential from OpenMM 8.1.141.
@@ 2.5 Analysis
Binding free energies and their corresponding uncertainties from the perturbation energy samples were computed using the Unbinned Weighted Histogram Analysis Method (UWHAM).42 The resulting relative binding free energies ($\Delta\Delta G$) were averaged over the three repeats and used as the predicted value, and the standard deviation across the repeats was used as an error estimate. Absolute $\Delta G$ values were computed via a maximum likelihood estimator with cinnabar for the obtained $\Delta\Delta G$ values.43 The accuracy of the methods was evaluated with several metrics, both from an error and ranking perspective. Relative and absolute binding free energies were compared to experimental measurements of mean absolute error (MAE), root mean square error (RMSE), and Kendall Tau correlation coefficient. Additionally, we evaluated the methods’ ability to prioritize the most active compounds in each dataset by analyzing top compound identification. Specifically, we examined whether the methods correctly identified the top 5 ligands and the top 30% of ligands ranked by experimental binding affinity. For the top 5 metrics, we assessed how many of the five most active compounds were included in the predicted top 5. For the top 30% metric, we determined the percentage of ligands within the experimental top 30% that were also present in the predicted top 30%. All results are reported with 95% confidence intervals, calculated using 1000 bootstrap samples. To ensure a fair comparison and minimize discrepancies that can appear from differences in algorithm implementation, we recalculated the metrics and converted $\Delta G$ values for OPLS4 with FEP+ using our analysis scripts. These calculations were performed by incorporating the $\Delta\Delta G$ values prior to the cycle closure correction into the algorithm. The initial structures and the results of all RBFE simulations are available on GitHub (refer to the Data Availability section for further details).
@@ 3.1 QuantumBind-RBFE calculations
Table S1 presents the RMSE, MAE, and Kendall tau correlation values for the evaluated targets, comparing GAFF2, OPLS4, and AceForce 1.0. These metrics provide insight into both the accuracy (via RMSE and MAE) and ranking performance (via Kendall tau) of each method across a diverse set of protein-ligand systems. $\Delta\Delta G$ values for the aforementioned systems and methods are included in the supporting information (Supporting TableS7 and Supporting FigureS1. Since this work uses a different charge model (RESP) compared to our previous publications (AM1BCC), we compared the results obtained here with those from earlier studies (Supporting Tables S3, S5, S4). For most targets, the results are very similar between the two charge models. Notably, in certain cases, the improvement with RESP charges is considerable. While the limitations and possible improvements of this approach are discussed in later sections, the overall improvement observed with RESP charges motivated their use for the QuantumBind-RBFE calculations presented in this work.
@@
QuantumBind-RBFE demonstrates improved performance compared to GAFF2 across most targets. When considering all computed data points, AceForce 1.0 achieves better overall metrics, with a notable decrease in RMSE from 1.17 kcal/mol for GAFF2 to 0.99 kcal/mol for AceForce, and a reduction in MAE from 0.90 to 0.79 kcal/mol. The Kendall tau correlation is slightly better for AceForce 1.0 (0.59 vs. 0.55 for GAFF2), though the difference is relatively modest. Importantly, AceForce consistently achieves a mean absolute error (MAE) below 1.0 kcal/mol for all targets, underscoring its robust accuracy across diverse systems.
@@
Despite the similar accuracy in terms of RMSE and MAE across many targets, AceForce 1.0 consistently outperforms GAFF2 in ranking capabilities. Kendall tau correlations are higher for all targets except MCL1 and THROMBIN. The lower correlation for THROMBIN can be attributed to the small size of the dataset (N = 11), where even a single outlier can disproportionately affect the correlation metric, as observed in previous studies.44 For MCL1, the reduced Kendall tau (0.28 for AceForce vs. 0.44 for GAFF2) is primarily due to the overprediction of binding affinities for five ligands in the series, which significantly impacts ranking performance.
@@
AceForce 1.0 demonstrates competitive performance compared to the reported results with OPLS4,28 though the latter generally achieves lower RMSE and MAE values across most targets. When considering all computed data points, OPLS4 outperforms AceForce 1.0 with a lower RMSE (0.78 vs. 0.99 kcal/mol) and MAE (0.61 vs. 0.79 kcal/mol). Kendall tau correlations are also slightly higher for OPLS4 (0.66 vs. 0.59 for AceForce), indicating better overall ranking accuracy.
@@
Overall, OPLS4 show still superior RMSE for most targets, while comparable ranking performance to AceForce 1.0 where some targets are better ranked by OPLS4, some by AceForce 1.0. It is worth noting that the methodologies for RBFE calculations differ between QuantumBind and FEP+, including distinct protocols and molecular dynamics engines. These differences can introduce additional variability in the results. The observed differences are also usually small and often within the margin of error, emphasizing the promise of NNPs as a viable alternative for RBFE calculations.
@@
Finally, we highlight the significant improvement AceForce 1.0 provides over GAFF2 when using the same protocol and molecular dynamics engine. This underscores the robustness of AceForce 1.0 as an advanced tool for RBFE predictions in drug discovery workflows. The parameterization of small molecules with NNPs, even using the NNP/MM scheme, becomes trivial, as it does not require the calculation and fitting of torsion scans, greatly simplifying the process.
@@
Figure 3: (Top) Root Mean Squared Error (RMSE) and (bottom) Kendall tau correlation for the $\Delta G$s of each protein-ligand system calculated in combination with different approaches: GAFF2 (teal), reported estimates using FEP+ with the OPLS4 forcefield(yellow) and AceForce 1.0 (blue).
@@
Figure 4: Scatterplots of predicted $\Delta G$ values for each evaluated system using AceForce 1.0. The green and yellow shaded areas represent absolute error thresholds of 1 kcal/mol and 2 kcal/mol, respectively. Additional metrics, including mean absolute error (MAE), root mean square error (RMSE), Kendall tau correlation ($\tau$), and top 30% and top 5 compound identification accuracy, are also displayed. 95% confidence interval values for the relevant metrics are shown in brackets.

### v2

@@ Abstract
Accurate prediction of protein-ligand binding affinities is crucial in drug discovery, particularly during hit-to-lead and lead optimization phases, however, limitations in ligand force fields continue to impact prediction accuracy. In this work, we validate relative binding free energy (RBFE) accuracy using neural network potentials (NNPs) for the ligands. We utilize a novel NNP model, AceFF 1.0, based on the TensorNet architecture for small molecules that broadens the applicability to diverse drug-like compounds, including all important chemical elements and supporting charged molecules. Using established benchmarks, we show overall improved accuracy and correlation in binding affinity predictions compared with GAFF2 for molecular mechanics and ANI2-x for NNPs. Slightly less accuracy but comparable correlations with OPLS4. We also show that we can run the NNP simulations at 2 fs timestep, at least two times larger than previous NNP models, providing significant speed gains. The results show promise for further evolutions of free energy calculations using NNPs while demonstrating its practical use already with the current generation. The code and NNP model are publicly available for research use.
@@ 1 Introduction
Recently, an alternative approach has emerged based on neural network potentials (NNPs) to use fast neural network function approximation of the quantum mechanic’s energy surface12. Based on these architectures, a few methods have been made available which we will consider in this work. MACE-based models13 are very accurate but computationally expensive. This is not a problem in material science where the alternative is very expensive quantum chemistry methods. For biology-oriented applications, ANI-models14 are much faster but at the expenses of accuracy, still providing a very good trade-off. AIMNet215 extends ANI with a different architecture to more elements and molecules, still adding on computational costs. The equivariant transform (ET) architecture16 and TensorNet17 aimed to provide state-of-the-art accuracy without sacrificing computational efficiency, being taylored specifically for drug discovery. These two are integrated in the TorchMD-Net software framework 18, 16, 19 and a new NNP model has been made available, AceFF20. Molecular dynamics (MD) simulations based on NNPs promise accurate simulations at reasonable costs compared with density functional theory (DFT) methods, but do represent an increase in computational costs compared to traditional molecular mechanics, yet their future is very promising 21. Machine-learned potentials employ relatively short cutoff distances (typically around 5 Å) for efficiency. However, message-passing architectures—such as those based on TensorNet—effectively extend the receptive field beyond the nominal cutoff.17 For example, a single interaction layer with a 5 Å cutoff can yield an effective range of approximately 10 Å. Moreover, in our NNP/MM setup, only the ligand’s intramolecular interactions are modeled by the NNP. Long-range electrostatic is taken care of by the standard classical potential at this time.
Recent studies by Karwounopoulos et al.22 have explored an alternative approach in which torsional profiles in classical force fields are fitted via machine learning as a means to approximate the benefits of using an NNP with mechanical embedding. Their results indicate that this end‐state corrected MM/ML method does not offer significant improvements over standard MM in capturing the energetic contributions of ligand strain. In contrast, NNP/MM approach directly employs a neural network potential to model the complete intramolecular energy surface of the ligand, providing a more comprehensive treatment of ligand strain and other subtle energetic effects.
@@
Recently, we explored the integration of NNPs into RBFE calculations with promising results.23 Specifically, we used ANI-2x14 with good results. However, ANI-2x cannot handle charged species and has a restricted range of supported atom elements. These limitations restrict the applicability across the diverse chemical space encountered in drug discovery. Recently, alternative approaches have emerged that address these shortcomings by integrating quantum-mechanical insights into the potential. Notably, the QDpi models24, 25, use a fast QM/MM-MLP framework to model both intra- and intermolecular interactions, including challenging cases such as tautomers and different protonation states. These models have demonstrated improved performance in reproducing properties like proton affinities. Furthermore, Crha et al. extended their QM/MM framework by integrating a machine‐learned potential to predict both the QM region and its fully polarized buffer’s energies, thereby enabling direct alchemical free‐energy perturbations as demonstrated in the methanol-to-methane transformation in water.26 In this work, we extend our previous work23 by testing QuantumBind-RBFE, our NNP/MM approach27 for RBFE calculations, using AceFF. The AceFF 1.020 model supports a broad range of atom elements, including charged molecules, and is tailored for accurate and efficient RBFE calculations across diverse molecular systems. To validate the performance of RBFE calculations using NNPs, we conducted a comprehensive benchmarking study using established datasets, including both charged and neutral ligands. Our results show that we can improve the accuracy of RBFE calculations compared to traditional force fields and our previous test with ANI-2x, reaching state-of-the-art correlations.
@@ 2.1 RBFE calculation within the NNP/MM scheme
Calculations were performed using an NNP/MM approach, which combines neural network potentials (NNP) for high-accuracy modeling of ligand interactions (e.g., capturing internal strain) and molecular mechanics (MM) for the remainder of the system, including all ligand-protein and ligand-solvent interactions.27 This hybrid method allows the ligand to be simulated with the more accurate neural network potential, while the surrounding protein environment is treated with classical molecular mechanics for computational efficiency.(Figure 1) In our NNP/MM scheme, the system is divided into NNP and MM regions, akin to the partitioning used in QM/MM simulations. The total potential energy ($V$) is calculated as:
@@
where $V_{\text{NNP}}$ describes the ligand’s intramolecular interactions using the NNP, and $V_{\text{MM}}$ accounts for the classical MM contributions of the protein and solvent. The coupling term $V_{\text{NNP-MM}}$, representing the nonbonded interactions (electrostatic and van der Waals) between the ligand and its environment is computed entirely using MM. This mechanical embedding scheme ensures that while the ligand–protein interactions are modeled with established MM terms, the ligand’s internal strain—which plays a critical role in binding free energy predictions—is captured at a higher level of theory via the NNP. Although “range corrected” neural network models that incorporate short-range interactions with the MM environment have shown promise in other contexts,28 our current implementation uses a purely mechanical embedding approach.27
@@
We utilized the Alchemical Transfer Method (ATM) to perform the RBFE calculations.29 This methodology has been previously validated across multiple benchmarks,30, 31 demonstrating its reliability for accurate binding free energy predictions. Notably, ATM was also employed in our recent work validating its use with NNPs, specifically ANI2x.23 In this study, we extend the previous work substantially.
@@ 2.2 Dataset selection
We selected a widely recognized public benchmark dataset that serves as a reference for RBFE studies. The BACE, CDK2, JNK1, MCL1, P38, THROMBIN, and TYK2 targets are part of Wang et al.’s dataset, commonly known as the ”JACS dataset” or ”Schrödinger dataset” 32. We had to omit PTP1B from the evaluation study since all ligands in the series have charge -2 and AceFF, for now, can only compute ligands with charges -1,0 and +1. Compared to our previous work, we can now perform the benchmark in completion. This dataset provides a total of 7 protein targets, 179 ligands, and 280 edges. Comparisons will be made with GAFF2 and FEP+ with the OPLS4 force field33.
@@ 2.3 NNP model
AceFF 1.0 is the first version of a new family of potentials 20. It uses TensorNet 1-layer 17 trained on Acellera34’s internal proprietary dataset of forces and energies computed the wB97M-V/def2-tzvppd level of theory. This work is the first test of this model for accuracy in RBFE calculations. A basic evaluation of the model on the Seller’s torsion scan 35 benchmark shows promising performance (Figure 2). For all molecules in the test, we performed the full torsion scans, minimizing the geometry with constrained torsion angles using the forces and energy from each calculation method. For comparison, we did the same procedure with ANI-2x14 from TorchANI36, xTB37, AIMNet215, and MACE-OFF13. We used GeomeTRIC38 for the constrained optimizations. The MAE between the curves for each molecule is shown in Figure 2. The AceFF model is among the best performing and comparable to MACE but at a much faster computational speed. AceFF is using a faster TensorNet 1-layer model, instead of the 2-layer model generally used in 17 for maximum accuracy, which has a speed of 80ns/day for a small molecule19 (timestep 1fs) compared to MACE-OFF23-Small with 7.5ns/day13.
Figure 2: Seller’s torsion scan benchmark, comparing torsion accuracy for a series of potentials against a CCSD(T)/CBS baseline.
@@ 2.4 QuantumBind-RBFE Calculation details
The workflow in this project follows the same procedure as in our previous work23. Ligands were parameterized with GAFF26, 7 with RESP charges, which are calculated with the OpenFF-Recharge package.39 The ligand topologies were generated using the parameterize40 tool. Starting from the protein and prepared ligand structures the protein-ligand complex systems are prepared with HTMD41. Following the protocol from previous works29, 30, 23 one of the ligands is displaced outside of the binding site via a displacement vector. We ensure that the ligand stays more than 15 Å away from the protein. Next the system is solvated with TIP3P waters with a padding of 10 Å and ions Na+ and Cl- are added to neutralize the system at a concentration of 0.15 M. By placing both ligands in the same simulation box with a dual topology approach42, we avoid issues related to charge-change modifications that can occur when RBFE calculations are performed in separate simulations. Finally, atom indexes for ligand alignment were performed, selecting three reference atoms for each ligand. For more information on the selection of these atoms please check previous publications.29, 30.
The energy minimization, thermalization, and equilibration steps followed the procedures described in our previous work.30. RBFE simulations were run in triplicate for each edge for an ensemble of 70 ns per replica, a similar amount of simulation time to our previous23 and similar43 studies. We used the Amber ff14SB parameters44, 45 as well as the TIP3P water model. In all simulations, bonds involving hydrogen atoms were constrained, which mitigates high-frequency vibrational motions and enables the use of larger timesteps. However, we also performed NVE simulations of systems without any constraints and verified that they can similarly conserve the energy as classical MD at the equivalent timestep. Details on the alchemical schedule can be found in Table S1. Classical RBFE simulations were run at a $4$ fs timestep while the NNP/MM runs were run at $2$ fs timestep with a Langevin thermostat for both cases. Previously, we already evaluated how the accuracy difference between 1fs and 4fs runs is minimal for MM runs.23 In the results section we discuss the reasoning behind selecting such a timestep for the NNP/MM runs. All calculations were run on GPUGRID.net46. The parallel replica exchange molecular dynamics simulations were conducted using QuantumBind-RBFE our implementation of ATM, HTMD41 and the ATMForce potential from OpenMM 8.1.147.
@@ 2.5 Analysis
Binding free energies and their corresponding uncertainties from the perturbation energy samples were computed using the Unbinned Weighted Histogram Analysis Method (UWHAM).48 The resulting relative binding free energies ($\Delta\Delta G$) were averaged over the three repeats and used as the predicted value, and the standard deviation across the repeats was used as an error estimate. Absolute $\Delta G$ values were computed via a maximum likelihood estimator with cinnabar for the obtained $\Delta\Delta G$ values.49 The accuracy of the methods was evaluated with several metrics, both from an error and ranking perspective. Relative and absolute binding free energies were compared to experimental measurements of mean absolute error (MAE), root mean square error (RMSE), and Kendall Tau correlation coefficient. Additionally, we evaluated the methods’ ability to prioritize the most active compounds in each dataset by analyzing top compound identification. Specifically, we examined whether the methods correctly identified the top 5 ligands and the top 30% of ligands ranked by experimental binding affinity. For the top 5 metrics, we assessed how many of the five most active compounds were included in the predicted top 5. For the top 30% metric, we determined the percentage of ligands within the experimental top 30% that were also present in the predicted top 30%. All results are reported with 95% confidence intervals, calculated using 1000 bootstrap samples. To ensure a fair comparison and minimize discrepancies that can appear from differences in algorithm implementation, we recalculated the metrics and converted $\Delta G$ values for OPLS4 with FEP+ using our analysis scripts. These calculations were performed by incorporating the $\Delta\Delta G$ values prior to the cycle closure correction into the algorithm. The initial structures and the results of all RBFE simulations are available on GitHub (refer to the Data Availability section for further details).
@@ 3.1 QuantumBind-RBFE calculations
Figure 3 and Tables S2,S3 presents the RMSE, MAE, and Kendall tau correlation values for the evaluated targets, comparing GAFF2, OPLS4, and AceFF 1.0. These metrics provide insight into both the accuracy (via RMSE and MAE) and ranking performance (via Kendall tau) of each method across a diverse set of protein-ligand systems. Scatter plots related to the AceFF 1.0 runs can be found in Figure 4. $\Delta\Delta G$ values for the aforementioned systems and methods are included in the supporting information (Table S4 and Figure S1. Since this work uses a different charge model (RESP) compared to our previous publications (AM1BCC), we compared the results obtained here with those from earlier studies (Tables S5-S7). For most targets, the results are very similar between the two charge models. Notably, in certain cases, the improvement with RESP charges is considerable. While the limitations and possible improvements of this approach are discussed in later sections, the overall improvement observed with RESP charges motivated their use for the QuantumBind-RBFE calculations presented in this work.
@@
QuantumBind-RBFE demonstrates improved performance compared to GAFF2 across most targets. When considering all computed data points, AceFF 1.0 achieves better overall metrics, with a notable decrease in RMSE from 1.17 kcal/mol for GAFF2 to 0.99 kcal/mol for AceFF, and a reduction in MAE from 0.90 to 0.79 kcal/mol. The Kendall tau correlation is slightly better for AceFF 1.0 (0.59 vs. 0.55 for GAFF2), though the difference is relatively modest. Importantly, AceFF consistently achieves a mean absolute error (MAE) below 1.0 kcal/mol for all targets, underscoring its robust accuracy across diverse systems.
@@
Despite the similar accuracy in terms of RMSE and MAE across many targets, AceFF 1.0 consistently outperforms GAFF2 in ranking capabilities. Kendall tau correlations are higher for all targets except MCL1 and THROMBIN. The lower correlation for THROMBIN can be attributed to the small size of the dataset (N = 11), where even a single outlier can disproportionately affect the correlation metric, as observed in previous studies.50 For MCL1, the reduced Kendall tau (0.28 for AceFF vs. 0.44 for GAFF2) is primarily due to the overprediction of binding affinities for five ligands in the series, which significantly impacts ranking performance.
We analyzed trajectories from both AceFF 1.0 and GAFF2 runs and found that the ligands explore similar conformational spaces in both cases. This indicates that the improved RBFE predictions with AceFF 1.0 mainly come from a better treatment of the ligand’s internal energetics,particularly internal strain, rather than from differences in overall conformational sampling.
@@
AceFF 1.0 demonstrates competitive performance compared to the reported results with OPLS4,33 though the latter generally achieves lower RMSE and MAE values across most targets. When considering all computed data points, OPLS4 outperforms AceFF 1.0 with a lower RMSE (0.78 vs. 0.99 kcal/mol) and MAE (0.61 vs. 0.79 kcal/mol). Kendall tau correlations are also slightly higher for OPLS4 (0.66 vs. 0.59 for AceFF), indicating better overall ranking accuracy.
@@
Overall, OPLS4 show still superior RMSE for most targets, while comparable ranking performance to AceFF 1.0 where some targets are better ranked by OPLS4, some by AceFF 1.0. It is worth noting that the methodologies for RBFE calculations differ between QuantumBind and FEP+, including distinct protocols and molecular dynamics engines. These differences can introduce additional variability in the results. The observed differences are also usually small and often within the margin of error, emphasizing the promise of NNPs as a viable alternative for RBFE calculations.
@@
Finally, we highlight the significant improvement AceFF 1.0 provides over GAFF2 when using the same protocol and molecular dynamics engine. This underscores the robustness of AceFF 1.0 as an advanced tool for RBFE predictions in drug discovery workflows. The parameterization of small molecules with NNPs, even using the NNP/MM scheme, becomes trivial, as it does not require the calculation and fitting of torsion scans, greatly simplifying the process.
@@
Figure 3: (Left) Root Mean Squared Error (RMSE) and (right) Kendall tau correlation for the $\Delta G$s of each protein-ligand system calculated in combination with different approaches: GAFF2 (teal), reported estimates using FEP+ with the OPLS4 forcefield(yellow) and AceFF 1.0 (blue).
@@
Figure 4: Scatterplots of predicted $\Delta G$ values for each evaluated system using AceFF 1.0. The green and yellow shaded areas represent absolute error thresholds of 1 kcal/mol and 2 kcal/mol, respectively. Additional metrics, including mean absolute error (MAE), root mean square error (RMSE), Kendall tau correlation ($\tau$), and top 30% and top 5 compound identification accuracy, are also displayed. 95% confidence interval values for the relevant metrics are shown in brackets.